divide 29 in to two parts so that the sum of the sqare of the parts is 425 .find the value of each parts.
Answer by Alan3354(69443)

divide 29 in to two parts so that the sum of the sqare of the parts is 425 .find the value of each parts.
-------------
x + y = 29
x^2 + y^2 = 425
------
x^2 + (29 - x)^2 = 425
2x^2 - 58x + 841 = 425
2x^2 - 58x + 416 = 0
x^2 - 29x + 208 = 0
etc